# Twisted character formula for toral supercuspidal representations

### v1

@@ 5.2. Twist and intertwiner
Then we get an intertwiner between $\operatorname{Ind}_{K^{\eta}}^{G}\rho_{(\mathbf{S},\vartheta)^{\eta}}$ and $\operatorname{Ind}_{K^{\eta}}^{G}\rho_{(\mathbf{S},\vartheta)}^{\eta}$ given by
$\operatorname{Ind}_{K^{\eta}}^{G}\rho_{(\mathbf{S},\vartheta)^{\eta}}\xrightarrow{\sim}\operatorname{Ind}_{K^{\eta}}^{G}\rho_{(\mathbf{S},\vartheta)}^{\eta}\colon f\mapsto I_{\rho_{(\mathbf{S},\vartheta)}}^{\eta}\circ f.$
@@ 6.1. Structure of the Heisenberg quotient
For $\Gamma\alpha\in\dot{\Phi}(\mathbf{G},\mathbf{S})$, we put
$\mathbf{V}_{\Gamma\alpha}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\bigoplus_{\beta\in\Gamma\alpha}\mathbf{V}_{\beta}\quad\text{and}\quad V_{\Gamma\alpha}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathbf{V}_{\Gamma\alpha}^{\Gamma}.$
Then, for any $\Gamma\alpha\in\dot{\Phi}(\mathbf{G},\mathbf{S})$, we have
$V_{\alpha}\xrightarrow{\sim}V_{\Gamma\alpha}=\Bigl(\bigoplus_{\beta\in\Gamma\alpha}\mathbf{V}_{\beta}\Bigr)^{\Gamma}\colon X_{\alpha}\mapsto\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}\sigma(X_{\alpha}).$
@@
(4) $V\cong\bigoplus_{\Gamma\alpha\in\dot{\Xi}}V_{\Gamma\alpha}=\bigoplus_{\Sigma\alpha\in\ddot{\Xi}}V_{\Sigma\alpha},$
where we put $V_{\Sigma\alpha}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V_{\Gamma\alpha}\oplus V_{-\Gamma\alpha}$ for $\alpha\in\Xi_{\mathrm{asym}}$ and $V_{\Sigma\alpha}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V_{\Gamma\alpha}$ for $\alpha\in\Xi_{\mathrm{sym}}$.
@@
where $c\in\mathbb{F}_{p}^{\times}$ is a constant determined by the fixed identification $\mu_{p}\cong\mathbb{F}_{p}$. Here, $\langle X^{\ast},[X_{1},X_{2}]\rangle\in k$ denotes the pairing of $X^{\ast}\in\mathfrak{s}^{\ast}_{r:r+}$ with the $\mathfrak{s}$-part of $X_{1},X_{2}]\in\mathfrak{g}_{\mathbf{x},r:r+}$ (i.e., the trivial isotypic component with respect to the $\mathbf{S}$-action). In fact, the above decomposition (4) gives an orthogonal decomposition of $V$ into symplectic subspaces. Each symplectic subspace $V_{\Sigma\alpha}$ is described as follows.
@@
Suppose that $\alpha\in\Xi_{\mathrm{asym}}$. We put $V_{\pm\alpha}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V_{\alpha}\oplus V_{-\alpha}$. Recall that the identification $V_{\alpha}\cong V_{\Gamma\alpha}\subset V$ is given by $X_{\alpha}\mapsto\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}\sigma(X_{\alpha})$. By noting that, for any $\alpha_{1},\alpha_{2}\in\Xi$, we have $\langle X^{\ast},[X_{\alpha_{1}},X_{\alpha_{2}}]\rangle\neq 0$ only if $\alpha_{2}=-\alpha_{1}$, the resulting symplectic form $V_{\pm\alpha}\times V_{\pm\alpha}\rightarrow\mathbb{F}_{p}$ maps $(X_{\alpha}+X_{-\alpha},Y_{\alpha}+Y_{-\alpha})$ to
@@ 6.2. Intertwiner of Heisenberg–Weil representations
Recall that we fixed a topologically semisimple element $\underline{\eta}\in\tilde{S}$ (Section 5.2). Hence any element $\eta^{\prime}\in\tilde{S}$ is written as $\eta^{\prime}=s\cdot\underline{\eta}$ with a unique element $s\in S$. Note that the action of $\underline{\eta}]$ on $\bm{\mathfrak{g}}$ induces an action on the set $\Phi(\mathbf{G},\mathbf{S})$ of order $l$, which does not depend on the choice of $\underline{\eta}\in\tilde{S}$. By abuse of notation, let us write $\theta_{\mathbf{S}}$ for this action and $\Theta_{\mathbf{S}}$ for the group $\langle\theta_{\mathbf{S}}\rangle$ generated by $\theta_{\mathbf{S}}$. To be more precise, for any $\alpha\in\Phi(\mathbf{G},\mathbf{S})$, $\theta_{\mathbf{S}}(\alpha)$ is the root given by $\theta_{\mathbf{S}}(\alpha)=\alpha\circ[\underline{\eta}]^{-1}$. Whenever there is no risk of confusion, we abbreviate $\theta_{\mathbf{S}}(\alpha)$ even as $\theta(\alpha)$. We note that, for $X_{\alpha}\in\bm{\mathfrak{g}}_{\alpha}$, $\underline{\eta}](X_{\alpha})$ belongs to $\mathfrak{g}_{\theta(\alpha)}$. We also note that, as $\underline{\eta}$ is $F$-rational, the actions of $\Theta_{\mathbf{S}}$ and $\Sigma=\Gamma\times\{\pm 1\}$ on $\Phi(\mathbf{G},\mathbf{S})$ commute. Especially, the symmetry of $\Phi(\mathbf{G},\mathbf{S})$ is preserved by $\Theta_{\mathbf{S}}$.
@@
Moreover, each $V_{\Sigma\alpha}$ is mapped onto $V_{\Sigma\theta(\alpha)}$, respectively. In particular, the action of $\Theta_{\mathbf{S}}$ on $\Phi(\mathbf{G},\mathbf{S})$ preserves $\Xi$.
@@
For $\alpha\in\Phi(\mathbf{G},\mathbf{S})$, let $m_{\alpha}$ be the order of $\Sigma\backslash(\Sigma\times\Theta_{\mathbf{S}})\alpha$. In other words, $m_{\alpha}$ is the smallest positive integer such that $\Sigma\theta^{m_{\alpha}}(\alpha)=\Sigma\alpha$ (hence $m_{\alpha}\mid l_{\alpha}$).
For $\Sigma\alpha\in\ddot{\Xi}$, let us write $(\omega_{\Sigma\alpha},W_{\Sigma\alpha})$ for the Heisenberg–Weil representation of $\operatorname{Sp}(V_{\Sigma\alpha})\ltimes\mathbb{H}(V_{\Sigma\alpha})$ with central character given by $\hat{\vartheta}$, which is unique up to isomorphism. Since the action $\underline{\eta}]$ on $V$ induces an symplectic isomorphism from $V_{\Sigma\alpha}$ to $V_{\Sigma\theta(\alpha)}$, an isomorphism from $\operatorname{Sp}(V_{\Sigma\alpha})\ltimes\mathbb{H}(V_{\Sigma\alpha})$ to $\operatorname{Sp}(V_{\Sigma\theta(\alpha)})\ltimes\mathbb{H}(V_{\Sigma\theta(\alpha)})$ is induced (for which we write $\underline{\eta}]_{\ast}$). Then the pull back $(\omega_{\Sigma\theta(\alpha)}^{\underline{\eta}},W_{\Sigma\theta(\alpha)})$ of the Heisenberg–Weil representation $W_{\Sigma\theta(\alpha)}$ of $\operatorname{Sp}(V_{\Sigma\theta(\alpha)})\ltimes\mathbb{H}(V_{\Sigma\theta(\alpha)})$ to $\operatorname{Sp}(V_{\Sigma\alpha})\ltimes\mathbb{H}(V_{\Sigma\alpha})$ via $\underline{\eta}]_{\ast}$ is isomorphic to $(\omega_{\Sigma\alpha},W_{\Sigma\alpha})$. For each $\Sigma\alpha\in\ddot{\Xi}$, we fix an intertwiner
$I_{\Sigma\alpha}^{\underline{\eta}}\colon\omega_{\Sigma\alpha}\xrightarrow{\sim}\omega_{\Sigma\theta(\alpha)}^{\underline{\eta}}.$
Recall that the representation $\omega=\omega_{(\mathbf{S},\vartheta)}$ is the Heisenberg–Weil representation of $\operatorname{Sp}(V)\ltimes\mathbb{H}(V)$ with central character $\hat{\vartheta}$. Since we have the decomposition (4), $\omega$ can be realized by tensoring Heisenberg–Weil representations $\omega_{\Sigma\alpha}$ for $\Sigma\alpha\in\ddot{\Xi}$ (see Section A.1). Furthermore, by tensoring the fixed intertwiners $I_{\Sigma\alpha}^{\underline{\eta}}$, we get an intertwiner between $\omega$ and its $\underline{\eta}]_{\ast}-$twist $\omega^{\underline{\eta}}$. Let us write $I_{\omega}^{\underline{\eta}}$ for the intertwiner obtained in this way:
@@
For each $\Sigma\alpha\in\ddot{\Xi}$, we put
$I^{\underline{\eta}}_{\Sigma\Theta(\alpha)}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}I_{\Sigma\theta^{m_{\alpha}-1}(\alpha)}^{\underline{\eta}}\circ\cdots\circ I_{\Sigma\alpha}^{\underline{\eta}}.$
@@
We still have not specified the choice of each $I^{\underline{\eta}}_{\Sigma\alpha}$ so far. This means that also $I_{\omega}^{\underline{\eta}}$ still has an ambiguity of a scalar multiple. Now we explain our choice of $I^{\underline{\eta}}_{\Sigma\alpha}$. For any $\Sigma\alpha\in\ddot{\Xi}$, note that $I^{\underline{\eta}}_{\Sigma\Theta(\alpha)}$ is an automorphism of $W_{\Sigma\alpha}$ such that
${W_{\Sigma\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{I^{\underline{\eta}}_{\Sigma\Theta(\alpha)}}$$\scriptstyle{\omega_{\Sigma\alpha}(g,h)}$${W_{\Sigma\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\omega_{\Sigma\alpha}([\underline{\eta}]^{m_{\alpha}}_{\ast}(g,h))}$${W_{\Sigma\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{I^{\underline{\eta}}_{\Sigma\Theta(\alpha)}}$${W_{\Sigma\alpha}}$
is commutative for any $(g,h)\in\operatorname{Sp}(V_{\Sigma\alpha})\ltimes\mathbb{H}(V_{\Sigma\alpha})$. Note that $\underline{\eta}]^{m_{\alpha}}$ is a symplectic automorphism of $V$ preserving $V_{\Sigma\alpha}$. Hence $I^{\underline{\eta}}_{\Sigma\Theta(\alpha)}$ must be a scalar multiple of the Heisenberg–Weil action $\omega_{\Sigma\alpha}([\underline{\eta}]^{m_{\alpha}})$. We choose $I^{\underline{\eta}}_{\Sigma\alpha}$ for $\Sigma\alpha\in\ddot{\Xi}$ so that we have
$I^{\underline{\eta}}_{\Sigma\Theta(\alpha)}=\omega_{\Sigma\alpha}([\underline{\eta}]^{m_{\alpha}}).$
@@
We eventually do not specify the choice of $I^{\underline{\eta}}_{\Sigma\alpha}$ in a unique way. Our normalization just requires that their composition $I^{\underline{\eta}}_{\Sigma\Theta(\alpha)}=I_{\Sigma\theta^{m_{\alpha}-1}(\alpha)}^{\underline{\eta}}\circ\cdots\circ I_{\Sigma\alpha}^{\underline{\eta}}$ exactly coincides with $\omega_{\Sigma\alpha}([\underline{\eta}]^{m_{\alpha}})$. In fact, this is enough for our purpose, i.e., what affects the value of the twisted character is only $I^{\underline{\eta}}_{\Sigma\Theta(\alpha)}$, not each individual $I^{\underline{\eta}}_{\Sigma\alpha}$, by Proposition 6.4.
@@
$\operatorname{tr}\bigl(\omega([s])\circ I_{\omega}^{\underline{\eta}}\bigr)=\prod_{\alpha\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\Sigma\alpha}}([\eta^{\prime}_{0}]^{m_{\alpha}}).$
@@
$\operatorname{tr}\bigl(\omega([s])\circ I_{\omega}^{\underline{\eta}}\bigr)=\prod_{\alpha\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\Sigma\alpha}}([\eta^{\prime}]^{m_{\alpha}})$
by Proposition 6.4. Noting that the topologically unipotent part $\eta^{\prime}_{+}$ acts on $V_{\Gamma\alpha}$ and $V_{\Sigma\alpha}$ trivially via conjugation, we get the assertion. ∎

### v2

@@ 5.2. Twist and intertwiner
Then we get an intertwiner between $\operatorname{c-Ind}_{K^{\eta}}^{G}\rho_{(\mathbf{S},\vartheta)^{\eta}}$ and $\operatorname{c-Ind}_{K^{\eta}}^{G}\rho_{(\mathbf{S},\vartheta)}^{\eta}$ given by
$\operatorname{c-Ind}_{K^{\eta}}^{G}\rho_{(\mathbf{S},\vartheta)^{\eta}}\xrightarrow{\sim}\operatorname{c-Ind}_{K^{\eta}}^{G}\rho_{(\mathbf{S},\vartheta)}^{\eta}\colon f\mapsto I_{\rho_{(\mathbf{S},\vartheta)}}^{\eta}\circ f.$
@@ 6.1. Structure of the Heisenberg quotient
For $\dot{\alpha}\in\dot{\Phi}(\mathbf{G},\mathbf{S})$, we put
$\mathbf{V}_{\dot{\alpha}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\bigoplus_{\beta\in\dot{\alpha}}\mathbf{V}_{\beta}\quad\text{and}\quad V_{\dot{\alpha}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathbf{V}_{\dot{\alpha}}^{\Gamma}.$
Then, for any $\dot{\alpha}\in\dot{\Phi}(\mathbf{G},\mathbf{S})$, we have
$V_{\alpha}\xrightarrow{\sim}V_{\dot{\alpha}}=\Bigl(\bigoplus_{\beta\in\dot{\alpha}}\mathbf{V}_{\beta}\Bigr)^{\Gamma}\colon X_{\alpha}\mapsto\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}\sigma(X_{\alpha}).$
@@
(4) $V\cong\bigoplus_{\dot{\alpha}\in\dot{\Xi}}V_{\dot{\alpha}}=\bigoplus_{\ddot{\alpha}\in\ddot{\Xi}}V_{\ddot{\alpha}},$
where we put $V_{\ddot{\alpha}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V_{\dot{\alpha}}\oplus V_{-\dot{\alpha}}$ for $\alpha\in\Xi_{\mathrm{asym}}$ and $V_{\ddot{\alpha}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V_{\dot{\alpha}}$ for $\alpha\in\Xi_{\mathrm{sym}}$.
@@
where $c\in\mathbb{F}_{p}^{\times}$ is a constant determined by the fixed identification $\mu_{p}\cong\mathbb{F}_{p}$. Here, $\langle X^{\ast},[X_{1},X_{2}]\rangle\in k$ denotes the pairing of $X^{\ast}\in\mathfrak{s}^{\ast}_{r:r+}$ with the $\mathfrak{s}$-part of $X_{1},X_{2}]\in\mathfrak{g}_{\mathbf{x},r:r+}$ (i.e., the trivial isotypic component with respect to the $\mathbf{S}$-action). In fact, the above decomposition (4) gives an orthogonal decomposition of $V$ into symplectic subspaces. Each symplectic subspace $V_{\ddot{\alpha}}$ is described as follows.
@@
Suppose that $\alpha\in\Xi_{\mathrm{asym}}$. We put $V_{\pm\alpha}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V_{\alpha}\oplus V_{-\alpha}$. Recall that the identification $V_{\alpha}\cong V_{\dot{\alpha}}\subset V$ is given by $X_{\alpha}\mapsto\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}\sigma(X_{\alpha})$. By noting that, for any $\alpha_{1},\alpha_{2}\in\Xi$, we have $\langle X^{\ast},[X_{\alpha_{1}},X_{\alpha_{2}}]\rangle\neq 0$ only if $\alpha_{2}=-\alpha_{1}$, the resulting symplectic form $V_{\pm\alpha}\times V_{\pm\alpha}\rightarrow\mathbb{F}_{p}$ maps $(X_{\alpha}+X_{-\alpha},Y_{\alpha}+Y_{-\alpha})$ to
@@ 6.2. Intertwiner of Heisenberg–Weil representations
Recall that we fixed a topologically semisimple element $\underline{\eta}\in\tilde{S}$ (Section 5.2). Hence any element $\eta^{\prime}\in\tilde{S}$ is written as $\eta^{\prime}=s\cdot\underline{\eta}$ with a unique element $s\in S$. Note that the action of $\underline{\eta}]$ on $\bm{\mathfrak{g}}$ induces an action on the set $\Phi(\mathbf{G},\mathbf{S})$ of order $l$, which does not depend on the choice of $\underline{\eta}\in\tilde{S}$. By abuse of notation, let us write $\theta_{\mathbf{S}}$ for this action and $\Theta_{\mathbf{S}}$ for the group $\langle\theta_{\mathbf{S}}\rangle$ generated by $\theta_{\mathbf{S}}$. To be more precise, for any $\alpha\in\Phi(\mathbf{G},\mathbf{S})$, $\theta_{\mathbf{S}}(\alpha)$ is the root given by $\theta_{\mathbf{S}}(\alpha)=\alpha\circ[\underline{\eta}]^{-1}$. Whenever there is no risk of confusion, we abbreviate $\theta_{\mathbf{S}}(\alpha)$ and $\Theta_{\mathbf{S}}(\alpha)$ even as $\theta(\alpha)$ and $\Theta(\alpha)$, respectively. We note that, for $X_{\alpha}\in\bm{\mathfrak{g}}_{\alpha}$, $\underline{\eta}](X_{\alpha})$ belongs to $\mathfrak{g}_{\theta(\alpha)}$. We also note that, as $\underline{\eta}$ is $F$-rational, the actions of $\Theta_{\mathbf{S}}$ and $\Sigma=\Gamma\times\{\pm 1\}$ on $\Phi(\mathbf{G},\mathbf{S})$ commute. Especially, the symmetry of $\Phi(\mathbf{G},\mathbf{S})$ is preserved by $\Theta_{\mathbf{S}}$.
@@
Moreover, each $V_{\ddot{\alpha}}$ is mapped onto $V_{\theta(\ddot{\alpha})}$, respectively. In particular, the action of $\Theta_{\mathbf{S}}$ on $\Phi(\mathbf{G},\mathbf{S})$ preserves $\Xi$.
@@
For $\alpha\in\Phi(\mathbf{G},\mathbf{S})$, let $m_{\alpha}$ be the order of $\Sigma\backslash(\Sigma\times\Theta_{\mathbf{S}})\alpha$. In other words, $m_{\alpha}$ is the smallest positive integer such that $\theta^{m_{\alpha}}(\ddot{\alpha})=\ddot{\alpha}$ (hence $m_{\alpha}\mid l_{\alpha}$).
For $\ddot{\alpha}\in\ddot{\Xi}$, let us write $(\omega_{\ddot{\alpha}},W_{\ddot{\alpha}})$ for the Heisenberg–Weil representation of $\operatorname{Sp}(V_{\ddot{\alpha}})\ltimes\mathbb{H}(V_{\ddot{\alpha}})$ with central character given by $\hat{\vartheta}$, which is unique up to isomorphism. Since the action $\underline{\eta}]$ on $V$ induces an symplectic isomorphism from $V_{\ddot{\alpha}}$ to $V_{\theta(\ddot{\alpha})}$, an isomorphism from $\operatorname{Sp}(V_{\ddot{\alpha}})\ltimes\mathbb{H}(V_{\ddot{\alpha}})$ to $\operatorname{Sp}(V_{\theta(\ddot{\alpha})})\ltimes\mathbb{H}(V_{\theta(\ddot{\alpha})})$ is induced (for which we write $\underline{\eta}]_{\ast}$). Then the pull back $(\omega_{\theta(\ddot{\alpha})}^{\underline{\eta}},W_{\theta(\ddot{\alpha})})$ of the Heisenberg–Weil representation $W_{\theta(\ddot{\alpha})}$ of $\operatorname{Sp}(V_{\theta(\ddot{\alpha})})\ltimes\mathbb{H}(V_{\theta(\ddot{\alpha})})$ to $\operatorname{Sp}(V_{\ddot{\alpha}})\ltimes\mathbb{H}(V_{\ddot{\alpha}})$ via $\underline{\eta}]_{\ast}$ is isomorphic to $(\omega_{\ddot{\alpha}},W_{\ddot{\alpha}})$. For each $\ddot{\alpha}\in\ddot{\Xi}$, we fix an intertwiner
$I_{\ddot{\alpha}}^{\underline{\eta}}\colon\omega_{\ddot{\alpha}}\xrightarrow{\sim}\omega_{\theta(\ddot{\alpha})}^{\underline{\eta}}.$
Recall that the representation $\omega=\omega_{(\mathbf{S},\vartheta)}$ is the Heisenberg–Weil representation of $\operatorname{Sp}(V)\ltimes\mathbb{H}(V)$ with central character $\hat{\vartheta}$. Since we have the decomposition (4), $\omega$ can be realized by tensoring Heisenberg–Weil representations $\omega_{\ddot{\alpha}}$ for $\ddot{\alpha}\in\ddot{\Xi}$ (see Section A.1). Furthermore, by tensoring the fixed intertwiners $I_{\ddot{\alpha}}^{\underline{\eta}}$, we get an intertwiner between $\omega$ and its $\underline{\eta}]_{\ast}-$twist $\omega^{\underline{\eta}}$. Let us write $I_{\omega}^{\underline{\eta}}$ for the intertwiner obtained in this way:
@@
For each $\ddot{\alpha}\in\ddot{\Xi}$, we put
$I^{\underline{\eta}}_{\Theta(\ddot{\alpha})}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}I_{\theta^{m_{\alpha}-1}(\ddot{\alpha})}^{\underline{\eta}}\circ\cdots\circ I_{\ddot{\alpha}}^{\underline{\eta}}.$
@@
We still have not specified the choice of each $I^{\underline{\eta}}_{\ddot{\alpha}}$ so far. This means that also $I_{\omega}^{\underline{\eta}}$ still has an ambiguity of a scalar multiple. Now we explain our choice of $I^{\underline{\eta}}_{\ddot{\alpha}}$. For any $\ddot{\alpha}\in\ddot{\Xi}$, note that $I^{\underline{\eta}}_{\Theta(\ddot{\alpha})}$ is an automorphism of $W_{\ddot{\alpha}}$ such that
${W_{\ddot{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{I^{\underline{\eta}}_{\Theta(\ddot{\alpha})}}$$\scriptstyle{\omega_{\ddot{\alpha}}(g,h)}$${W_{\ddot{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\omega_{\ddot{\alpha}}([\underline{\eta}]^{m_{\alpha}}_{\ast}(g,h))}$${W_{\ddot{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{I^{\underline{\eta}}_{\Theta(\ddot{\alpha})}}$${W_{\ddot{\alpha}}}$
is commutative for any $(g,h)\in\operatorname{Sp}(V_{\ddot{\alpha}})\ltimes\mathbb{H}(V_{\ddot{\alpha}})$. Note that $\underline{\eta}]^{m_{\alpha}}$ is a symplectic automorphism of $V$ preserving $V_{\ddot{\alpha}}$. Hence $I^{\underline{\eta}}_{\Theta(\ddot{\alpha})}$ must be a scalar multiple of the Heisenberg–Weil action $\omega_{\ddot{\alpha}}([\underline{\eta}]^{m_{\alpha}})$. We choose $I^{\underline{\eta}}_{\ddot{\alpha}}$ for $\ddot{\alpha}\in\ddot{\Xi}$ so that we have
$I^{\underline{\eta}}_{\Theta(\ddot{\alpha})}=\omega_{\ddot{\alpha}}([\underline{\eta}]^{m_{\alpha}}).$
@@
We eventually do not specify the choice of $I^{\underline{\eta}}_{\ddot{\alpha}}$ in a unique way. Our normalization just requires that their composition $I^{\underline{\eta}}_{\Theta(\ddot{\alpha})}=I_{\theta^{m_{\alpha}-1}(\ddot{\alpha})}^{\underline{\eta}}\circ\cdots\circ I_{\ddot{\alpha}}^{\underline{\eta}}$ exactly coincides with $\omega_{\ddot{\alpha}}([\underline{\eta}]^{m_{\alpha}})$. In fact, this is enough for our purpose, i.e., what affects the value of the twisted character is only $I^{\underline{\eta}}_{\Theta(\ddot{\alpha})}$, not each individual $I^{\underline{\eta}}_{\ddot{\alpha}}$, by Proposition 6.4.
@@
$\operatorname{tr}\bigl(\omega([s])\circ I_{\omega}^{\underline{\eta}}\bigr)=\prod_{\ddot{\alpha}\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\ddot{\alpha}}}([\eta^{\prime}_{0}]^{m_{\alpha}}).$
@@
$\operatorname{tr}\bigl(\omega([s])\circ I_{\omega}^{\underline{\eta}}\bigr)=\prod_{\ddot{\alpha}\in\Theta_{\mathbf{S}}\backslash\ddot{\Xi}}\Theta_{\omega_{\ddot{\alpha}}}([\eta^{\prime}]^{m_{\alpha}})$
by Proposition 6.4. Noting that the topologically unipotent part $\eta^{\prime}_{+}$ acts on $V_{\dot{\alpha}}$ and $V_{\ddot{\alpha}}$ trivially via conjugation, we get the assertion. ∎